Calculate the energy required in joules to raise the temperature of 450 grams of water from 15°C to 85°C. (The specific heat capacity of water is 4.18 J/g/°C)
1
Expert's answer
2019-10-17T11:06:43-0400

The requaierd energy can be caclulated as


"Q = cm\Delta T"

Where "c" - the specific heat capacity of water, "m" - mass of water, "\Delta T" - temperature defference. We get


"Q = 4.18[{{\rm{J}} \over {{\rm{g}} \cdot {\rm{^\circ C}}}}] \cdot 450[{\rm{g}}] \cdot (85[{\rm{^\circ C}}] - 15[{\rm{^\circ C}}]) = 131670[{\rm{J}}]"